### v1

# IPPP-25-XXX

Abstract
I review the standard instanton framework for determining the $\theta$-dependence of instanton-dominated correlation functions in QCD. I then contrast these well-established semiclassical results with the recent assertion of Refs. [Ai:2020ptm, Ai:2024cnp], that $\theta$-phases are absent and that strong interactions preserve CP.
Dilute Instanton Gas approximation.
Consider QCD with massive quarks and include the conventionally normalised $\theta$-term, $\frac{i\theta}{16\pi^{2}}\,{\rm tr}\,F_{\mu\nu}\tilde{F}_{\mu\nu}$, in the Euclidean Lagrangian. Assume that the theory is weakly coupled so that a semiclassical treatment is reliable. Our goal is to compute the partition function $Z(\theta)$ and thereby determine the vacuum energy density as a function of $\theta$,
$E(\theta)=-\frac{\log Z(\theta)}{V_{4}},$ (1)
where $V_{4}$ is the four-dimensional Euclidean volume. To proceed, we further assume that the dominant contributions to the path integral for $Z(\theta)$ arise from expanding around classical field configurations consisting of $n+\bar{n}$ widely separated, non-interacting instantons and anti-instantons. Summing over all possible numbers $n$ of instantons and $\bar{n}$ of anti-instantons leads to the standard dilute instanton gas approximation (DIGA) formula [Callan:1976je, Coleman1985Aspects],
$Z(\theta)=\,\sum_{n=0}^{\infty}\sum_{\bar{n}=0}^{\infty}\frac{1}{n!\,\bar{n}!}\left(\int d^{4}x_{0}\,K_{I}\,e^{-\frac{8\pi^{2}}{g^{2}}+i\theta}\right)^{n}\left(\int d^{4}x_{0}\,K_{\bar{I}}\,e^{-\frac{8\pi^{2}}{g^{2}}-i\theta}\right)^{\bar{n}}.$ (2)
[…]
$K_{I}\,=\,|K|\,e^{i\alpha}\,,\quad K_{\bar{I}}\,=\,|K|\,e^{-i\alpha}\,,\quad\alpha\,\equiv\,\log{\rm arg}\det M_{f}$ (3)
where $|K|$ is a positive constant of mass-dimension 4, so that $\int d^{4}x_{0}\,|K|=V_{4}\,|K|$ is dimensionless. It is also convenient to define the combined phase parameter $\bar{\theta}\,=\,\theta+\alpha$.
The standard (and ultimately correct) procedure within the DIGA framework is to evaluate the infinite double sum on the right-hand side of (2) at finite Euclidean volume $V_{4}$,
$Z(\theta)=\,\sum_{n=0}^{\infty}\frac{1}{n!}\left(V_{4}\,|K|\,e^{-\frac{8\pi^{2}}{g^{2}}+i\bar{\theta}}\right)^{n}\times\,\sum_{\bar{n}=0}^{\infty}\frac{1}{\bar{n}!}\left(V_{4}\,|K|\,e^{-\frac{8\pi^{2}}{g^{2}}-i\bar{\theta}}\right)^{\bar{n}}=\,e^{\,2V_{4}|K|e^{-8\pi^{2}/g^{2}}\cos\bar{\theta}}$ (4)
and then divide by the volume the logarithm of the result, as in (1),
$E(\theta)=\,-\,\frac{\log Z(\theta)}{V_{4}}\,=\,-2|K|\,e^{-8\pi^{2}/g^{2}}\,\cos\bar{\theta}.$ (5)
For non-trivial values of $\bar{\theta}$, CP symmetry is broken by strong interactions. A popular solution to the resulting strong CP problem is provided by promoting the CP-violating phase $\bar{\theta}$ to a dynamical axion field $a(x)$ via $\bar{\theta}(x)=a(x)/f_{a}$, where $f_{x}$ is the axion decay constant. The axion potential is then read from (5),
$V(a)=\,2|K|\,e^{-8\pi^{2}/g^{2}}(1-\cos(a(x)/f_{a})).$ (6)
We emphasise that although the dilute instanton gas expression for $Z(\theta)$ involves a summation over all possible numbers of instantons and anti-instantons, the expression for $\log Z(\theta)$ is determined solely by the contribution of a single instanton together with its Hermitian conjugate – namely, a single anti-instanton. In diagrammatic language, this corresponds to the sum of two minimal connected diagrams: the single-instanton vertex and the single-anti-instanton vertex. These diagrams are ‘connected’ in the sense that a single instanton is an elementary semiclassical configuration whose vertex cannot be factorised into products of more basic ones.
The DIGA partition function $Z(\theta)=e^{-V_{4}\,E(\theta)}$ in (4) does receive contributions from multi-instanton–multi-anti-instanton sectors, but these enter only through disconnected diagrams. This is exactly as expected: the exponential of the sum of connected diagrams automatically generates all disconnected diagrams built from them.
  1. Note that, in principle, the dilute gas approximation can be systematically improved by incorporating 2-particle, 3-particle, and all higher-particle interactions. In statistical mechanics this procedure is known as the virial (or Mayer cluster) expansion. For example, the pressure $p$ of a non-dilute gas is
    $p=k_{B}T\,\frac{1}{V}\log Z(T,V,\mu),$ (7)
    where $Z(T,V,\mu)$ is the grand canonical partition function and $V$ is the volume. The quantity analogous to the vacuum energy $E(\theta)\,V_{4}$ in (5) is therefore $p\,V/(k_{B}T)$. The virial expansion gives
    $\frac{pV}{k_{B}T}=\log Z(T,V,\mu)=V\kappa\left(1+\kappa B_{2}(T)+\kappa^{2}B_{3}(T)+\ldots\right),$ (8)
    where $\kappa$ is the particle density (the analogue of the instanton density $|K|\,e^{-8\pi^{2}/g^{2}}$ in (5)), and $B_{2}$, $B_{3}$, etc. are the 2-particle, 3-particle, and higher-particle interaction coefficients that encode systematic corrections to the ideal gas approximation.
    Following the same logic, the expression (5) for the vacuum energy $E(\theta)$ will receive power corrections of the form $\sum_{k=2}^{\infty}B_{k}\left(|K|\,e^{-8\pi^{2}/g^{2}}\right)^{k}$ multiplied by trigonometric functions of integer multiples of $\bar{\theta}$. For each $k$ there is a $k$-cluster of interacting instanton and anti-instanton configurations; each such cluster corresponds to a connected diagram.
Key point:
Using the dilute instanton gas formalism to compute the partition function – and, more generally, any correlator – automatically generates the full set of disconnected instanton diagrams. The physically relevant connected part of this set consists solely of the single-instanton ($1$) and single-anti-instanton ($\bar{1}$) contributions, [^2] This statement applies specifically to the DIGA framework, where the instanton gas is taken to be ideal, i.e. all interaction terms $B_{k\geq 2}$ are set to zero by construction.
$E$ $=$ $E_{1}+E_{\bar{1}}\,=\,\frac{-\,\log Z_{\rm DIGA}}{V_{4}},$ (9)
$\langle{\rm Correlator}\rangle_{\rm conn}$ $=$ $\langle{\rm Correlator}\rangle_{1}+\langle{\rm Correlator}\rangle_{\bar{1}}\,=\,\frac{\langle{\rm Correlator}\rangle_{\rm DIGA}}{Z_{\rm DIGA}}.$ (10)
The single-instanton contributions, $E_{1}(\theta)$ and $\langle{\rm Correlator}\rangle_{1}(\theta)$, are entirely sufficient to determine the full $\theta$-dependence of physical observables. [^3] On the right-hand side of (10) one may also include the 0-instanton (perturbative) contribution, $\langle{\rm Correlator}\rangle_{0}$, if it is non-zero. This does not modify the $\theta$-dependence. These quantities can be computed directly simply by evaluating the one-instanton contributions, without any reference to or need for performing the infinite summations over $n$ and $\bar{n}$ appearing in the disconnected DIGA expressions on the right-hand sides of (9)-(10).
Let us now examine more closely the structure of a generic connected correlator in an instanton background, starting from the expression on the right-hand side of (10). We have
$\langle{\rm Correlator}\rangle_{\rm conn}\,=\,\frac{\langle{\rm Correlator}\rangle_{\rm DIGA}}{Z_{\rm DIGA}}\,=\,\frac{\sum_{n=0}^{\infty}\sum_{\bar{n}=0}^{\infty}\,\langle{\rm Correlator}\rangle_{n,\bar{n}}}{\sum_{n=0}^{\infty}\sum_{\bar{n}=0}^{\infty}\,Z_{n,\bar{n}}},$ (11)
where $\langle{\rm Correlator}\rangle_{n,\bar{n}}$ and $Z_{n,\bar{n}}$ denote contributions of the ideal gas involving $n$-instantons and $\bar{n}$-anti-instantons. These are given as follows,
$Z_{n,\bar{n}}$ $=$ $\frac{(V_{4}|E_{1}|e^{i\bar{\theta}})^{n}}{n!}\,\,\frac{(V_{4}|E_{1}|e^{-i\bar{\theta}})^{\bar{n}}}{\bar{n}!},$ (12)
$\langle{\rm Correlator}\rangle_{n,\bar{n}}$ $=$ $\langle{\rm Correlator}\rangle_{1}\,Z_{n-1,\bar{n}}\,+\,\langle{\rm Correlator}\rangle_{\bar{1}}\,Z_{n,\bar{n}-1}\,+\,\langle{\rm Correlator}\rangle_{0}\,Z_{n,\bar{n}}.$ (13)
We see that each of these expressions corresponds to an $(n+\bar{n})$-fold product of elementary single-instanton and single-anti-instanton contributions. For example, the expression on the right-hand side of (12) is simply the product of $n$ single-instanton factors and $\bar{n}$ single-anti-instanton factors (divided by the appropriate Bose–Einstein symmetry factor). Likewise, the first term on the right-hand side of (13) represents the one-instanton contribution to the correlator multiplied by (n1)(n-1) disconnected single-instanton contributions and by $\bar{n}$ single-anti-instanton contributions to the partition function. All such terms are disconnected diagrams built from $(n+\bar{n})$ instanton ‘blobs’, exactly as expected in the dilute instanton gas picture.
After carrying out the sums over $n$ and $\bar{n}$ in both the numerator and the denominator in (11), all of the $Z$-factors cancel, and we obtain the simple result,
$\frac{\langle{\rm Correlator}\rangle_{\rm DIGA}}{Z_{\rm DIGA}}\,=\,\langle{\rm Correlator}\rangle_{1}\,+\,\langle{\rm Correlator}\rangle_{\bar{1}}\,+\,\langle{\rm Correlator}\rangle_{0},$ (14)
which is the correct expression for the connected correlator $\langle{\rm Correlator}\rangle_{\rm conn}$ and is fully consistent with (10).
Comments:
  1. 1.
    The correlators in (14) do depend on $\theta$ and this dependence can be made explicit,
    $\frac{\langle{\rm Correlator}\rangle_{\rm DIGA}}{Z_{\rm DIGA}}\,=\,e^{i\bar{\theta}}\,\langle{\rm Correlator}\rangle_{1,\,\bar{\theta}=0}\,+\,e^{-i\bar{\theta}}\,\langle{\rm Correlator}\rangle_{\bar{1},\,\bar{\theta}=0}\,+\,\langle{\rm Correlator}\rangle_{\rm pert.th.}$ (15)
  2. 2.
    At no stage in deriving (14) we had to take an infinite volume limit.
  3. 3.
    The use of the dilute-instanton-gas formulae and the associated sums over instanton numbers is entirely unnecessary for evaluating $\langle{\rm Correlator}\rangle_{\rm conn}$. This quantity can be computed directly in the single-instanton approximation (with perturbative contributions added if desired), and given by the right-hand side of (14). Note that it is always the connected part of the correlator that is directly related to terms in the effective action.
  4. 4.
    When interactions are included, additional genuine multi-instanton contributions can appear in (14). These correspond to true multi-instanton ADHM solutions, whose effects have been explicitly computed in $N=2$ and $N=4$ SYM theories in Refs. [Dorey:1996hu, Nekrasov:2002qd, Dorey:1999pd]. Such multi-instanton contributions are unrelated to the dilute gas approximation; rather, they represent systematic $k$-instanton corrections in the sense of the virial expansion (8). These effects do not play a role in our present analysis of $\theta$-dependence to leading order in the instanton number.
Now we are ready to compare the outlined above standard calculation of $\theta$-dependence to an alternative approach advocated in Refs. [Ai:2020ptm, Ai:2024cnp].
The ‘absence of strong CP’ approach:
The authors of [Ai:2020ptm, Ai:2024cnp] proposed to re-organise the infinite two-fold summation in the DIGA formula (11) into two steps, and apply the infinite volume limit after completing the first sum but before the second.
Specifically, the sum over $n$ and $\bar{n}$ is re-arranged by first holding the net instanton number $\Delta n=n-\bar{n}$ fixed while summing over individual values of $n$, and second, summing over all values of $\Delta n$. In so far this procedure is harmless [^4] The radius of convergence of the exponent and also of the Bessel functions appearing below is infinite, so the convergence criterium for the sums is $V_{4}e^{-8\pi^{2}/g^{2}}|K|<\infty$ which works for any arbitrarily large finite volume. Equation (11) can be presented in an equivalent form as,
$\langle{\rm Correlator}\rangle_{\rm conn}\,=\,\frac{\langle{\rm Correlator}\rangle_{\rm DIGA}}{Z_{\rm DIGA}}\,=\,\,\lim_{N\to\infty}\frac{\sum_{\Delta n=-N}^{N}\sum_{n}\,\langle{\rm Correlator}\rangle_{n,\,\bar{n}=n+\Delta n}}{\sum_{\Delta n=-N}^{N}\sum_{n}\,Z_{n,\bar{n}}}.$ (16)
Performing the inner summations over $n$ at fixed $\Delta n$ gives,
$e^{i\Delta n\bar{\theta}}\left(I_{\Delta n+1}(x)\,\langle{\rm Correlator}\rangle_{1,\,\bar{\theta}=0}\,+\,I_{\Delta n-1}(x)\,\langle{\rm Correlator}\rangle_{\bar{1},\,\bar{\theta}=0}\,+\,I_{\Delta n}(x)\,\langle{\rm Correlator}\rangle_{0}\right),$ (17)
$e^{i\Delta n\bar{\theta}}\,I_{\Delta n}(x),$ (18)
for the expressions in the numerator and the denominator, respectively. Here $I_{\Delta n\pm 1}(x)$ and $I_{\Delta n}(x)$ are the modified Bessel functions and their argument is $x=2V_{4}\,e^{-8\pi^{2}/g^{2}}|K|$.
If we sum these expressions over $\Delta n$ and evaluate the ratio on the right-hand side of (16), we recover the result (14) for the connected correlator $\langle{\rm Correlator}\rangle_{\rm conn}$.
The approach advocated in Refs. [Ai:2020ptm, Ai:2024cnp] is different. These authors insist on taking the infinite volume limit before the infinite sum over $\Delta n$. Specifically, they define the correlator,
$\langle{\rm Correlator}\rangle_{\rm ACGT}\,=\,\lim_{N\to\infty}\,\lim_{V_{4}\to\infty}\frac{\sum_{\Delta n=-N}^{N}\sum_{n}\,\langle{\rm Correlator}\rangle_{n,\,\bar{n}=n+\Delta n}}{\sum_{\Delta n=-N}^{N}\sum_{n}\,Z_{n,\bar{n}}}$ (19)
Mathematically, there is no need to impose the infinite-volume limit either before or after performing the infinite sum: as already noted, the sums in (4) and (11) converge for any finite value of $x$, and moreover the dependence on the spacetime volume cancels in the ratio appearing in (15). Nevertheless, the authors of Refs. [Ai:2020ptm, Ai:2024cnp] were strongly drawn to the idea of imposing the limit $\lim_{V_{4}\to\infty}$, and do it prior to extending the summation over $\Delta n$ to an infinite domain – as prescribed in (19). In these settings, one makes use of the asymptotic values of the Bessel functions,
$I_{m}(x)\,\approx\,\frac{1}{\sqrt{2\pi x}}\,\,e^{x}\,,\quad{\rm as}\,\,x\to\infty,\quad{\rm for\,\,any}\,\,m\in Z,$ (20)
so that all ratios of $I_{\Delta n\pm 1}(x)/I_{m}(x)$ and $I_{\Delta n}(x)/I_{m}(x)$ are =1=1 in the $V_{4}\to\infty$ limit. Normalising the numerator and denominator expressions (17)-(18) by the same $I_{m}(x)$ and taking the limit $x\to\infty$, one obtains for the numerator of (19):
$\frac{1}{\sqrt{2\pi x}}\,\,e^{x}\left(\langle{\rm Correlator}\rangle_{1,\,\bar{\theta}=0}\,+\,\langle{\rm Correlator}\rangle_{\bar{1},\,\bar{\theta}=0}\,+\,\langle{\rm Correlator}\rangle_{0}\right)\left(\sum_{\Delta n=-N}^{N}e^{i\Delta n\bar{\theta}}\right),$ (21)
and for the denominator:
$\frac{1}{\sqrt{2\pi x}}\,\,e^{x}\left(\sum_{\Delta n=-N}^{N}e^{i\Delta n\bar{\theta}}\right),$ (22)
Taking the ratio of these expressions, one immediately sees that the entire $\theta$-dependence cancels out in (19),
$\langle{\rm Correlator}\rangle_{\rm ACGT}\,=\,\langle{\rm Correlator}\rangle_{1,\,\bar{\theta}=0}\,+\,\langle{\rm Correlator}\rangle_{\bar{1},\,\bar{\theta}=0}\,+\,\langle{\rm Correlator}\rangle_{\rm pert.th.}$ (23)
The authors of Refs. [Ai:2020ptm, Ai:2024cnp] take this cancellation to imply that physical observables, as defined by the order of limits in (19), are insensitive to the CP-violating parameter $\theta$. This assertion forms the core of their claim that CP violation is absent in QCD.
I will now argue that taking the infinite volume limit before completing the second infinite sum (i.e before taking $N\to\infty$ in (19)) invalidates the structure of the numerator and denominator in this expression as sums over disconnected diagrams.
Let us examine the denominator first. Equation (22) shows that the partition function takes the form,
$Z\,=\,\lim_{N\to\infty}\,\lim_{V_{4}\to\infty}\sum_{\Delta n=-N}^{N}\sum_{n}\,Z_{n,\bar{n}}\,=\,\lim_{N\to\infty}\,\left(\sum_{\Delta n=-N}^{N}e^{i\Delta n\bar{\theta}}\right)\,\lim_{x\to\infty}\,\frac{1}{\sqrt{2\pi x}}\,\,e^{x}$ (24)
To simplify matters further we can temporarily set $\bar{\theta}=0$, so that,
$Z\,=\,\left[\lim_{N\to\infty}(2N+1)\right]\times\left[\lim_{x\to\infty}\,\frac{1}{\sqrt{2\pi x}}\,\,e^{x}\right]$ (25)
Only the factor $e^{x}=e^{\,2V_{4}|K|e^{-8\pi^{2}/g^{2}}}$ in this expression can be interpreted as the sum of disconnected instanton and anti-instanton contributions, in accordance with the result (4) evaluated at $\bar{\theta}=0$. However, this interpretation is spoiled by the additional multiplicative factor
$\frac{1}{\sqrt{2\pi x}}\,=\,\frac{1}{\sqrt{2\pi}}\,\frac{e^{4\pi^{2}/g^{2}}}{\sqrt{2V_{4}|K|}},$
which introduces a square-root dependence on the instanton density together with an explicit inverse power of the (infinite) spacetime volume. This is further compounded by the overall factor $2N$ which itself diverges in the limit of $N\to\infty$.
Applying the same line of reasoning to the numerator, one finds that the expression (21) in (19) similarly lacks the structure of a sum over disconnected contributions to the correlator.
The bottom line is that the object formally defined in (19) and computed in (23) does not appear to possess any clear physical meaning or interpretation.
  1. 1.
    Since the framework of Refs. [Ai:2020ptm, Ai:2024cnp] does not attempt to go beyond the ideal, non-interacting instanton gas and does not incorporate genuine multi-instanton contributions, it is particularly puzzling that the instanton correlators arising in their construction differ from the leading, single-instanton contributions on the right-hand side of (14).
  2. 2.
    The role normally played by dressing the minimal single-instanton correlators with the dilute instanton gas is to generate all disconnected diagrams. Once the connected graphs are extracted from this enlarged set, one should recover precisely the original single-instanton correlators. The order-of-limits prescription imposed in (19) obscures this standard construction. Consequently, after taking these limits, the quantity $\langle{\rm Correlator}\rangle_{\rm ACGT}$ can no longer be interpreted as the proper connected part of the correlator within the DIGA approximation presented in (11).
  3. 3.
    The rationale in Refs. [Ai:2020ptm, Ai:2024cnp] for taking the large-$V_{4}$ limit before summing over all topological sectors $\Delta n$ was to localise instantons well inside $V_{4}$, sufficiently far from the boundary at infinity. This was intended to ensure that, in every topological sector, the gauge fields approach the required pure-gauge boundary conditions.
    However, such concerns can be safely bypassed by expanding path integral around BPST instantons in singular gauge. In this gauge, the instanton gauge fields fall off as $\rho^{2}/x^{3}$ and therefore approach zero – not a non-trivial pure gauge – at large $|x|$. The topological winding is instead encoded around the (pure-gauge) singularity at the instanton centre. The singularity itself is harmless, being a removable gauge artefact. Indeed, it is well known that instantons in singular (rather than regular) gauge are the correct configurations to use when computing, for example, scattering amplitudes: their rapid fall-off guarantees that the LSZ reduction of the instanton background is non-vanishing on-shell,
    $A^{a\,{\rm inst}}_{LSZ}(p,\lambda)\,=\,\lim_{p^{2}\to 0}p^{2}\epsilon^{\mu}(p,\lambda)\,A_{\mu}^{a\,{\rm inst}}(p)\,=\,\epsilon^{\mu}(p,\lambda)\,\bar{\eta}^{a}_{\mu\nu}p_{\nu}\,\frac{4i\pi^{2}\rho^{2}}{g}\,e^{ip\cdot x_{0}}\,\neq 0\,.$ (26)
    while regular-gauge instantons would give a trivially vanishing contribution.
    Overall, I see no compelling reason to impose the infinite-volume limit in (19) in the first place.
  4. 4.
    In summary, we conclude that one should simply employ standard instanton calculus and compute the instanton contributions to the relevant connected correlators directly, as in (10):
    $\langle{\rm Correlator}\rangle_{\rm conn}\,=\,\langle{\rm Correlator}\rangle_{1}+\langle{\rm Correlator}\rangle_{\bar{1}}\,+\,\ldots$ (27)
    where the ellipsis denotes higher-order contributions from connected multi-instanton clusters. The correlators on the right-hand side exhibit the standard dependence on the $\theta$-parameter, as shown in (15), with $\langle{\rm Correlator}\rangle_{1}\propto e^{i\bar{\theta}}$ and $\langle{\rm Correlator}\rangle_{\bar{1}}\propto e^{-i\bar{\theta}}$.
  5. 5.
    Beyond QCD and strong CP: Instanton contributions in 4-dimensional gauge theories always carry complex phases which are naturally organised as a Fourier series in $q=e^{2\pi i\tau}$, where $\tau=\frac{\theta}{2\pi}+\frac{4\pi i}{g^{2}}$. In $\mathcal{N}=2$ Seiberg–Witten theory [Seiberg:1994rs] the low-energy effective action is encoded in the holomorphic prepotential $\mathcal{F}(a,\tau)$. For large $a$ one obtains an expansion
    $\mathcal{F}(a,\tau)=\mathcal{F}_{\mathrm{class+1loop}}(a,\tau)+\sum_{k\geq 1}\mathcal{F}_{k}(a)\,e^{2\pi ik\tau}.$ (28)
    where the summation is over all multi-instanton numbers $k$. The complex phases $e^{2\pi ik\tau}$ of instanton contributions [Dorey:1996hu, Nekrasov:2002qd] are essential to ensure holomorphy and correct monodromies of the prepotential [Seiberg:1994rs].
    In $\mathcal{N}=4$ SYM the factors of $e^{2\pi ik\tau}$ allow non-perturbative quantities to be organised into genuine modular or quasi-modular forms [Banks:1998nr, Green:1997tv] and [Dorey:1999pd]. They implement S-duality by providing the correct transformation behaviour under $SL(2,\mathbb{Z})$.
In this note I have focused exclusively on the instanton-based derivation of the $\theta$-dependence, leaving aside other complementary formulations and methods. For a recent ongoing discussion of these broader aspects, see Refs. [Benabou:2025viy, Ai:2025quf].
Acknowledgements
I thank Steve Abel, Rodrigo Alonso and Michael Spannowsky for useful discussions.